# Wild2Avatar: Rendering Humans Behind Occlusions

### v1

@@ Abstract
Rendering the visual appearance of moving humans from occluded monocular videos is a challenging task. Most existing research renders 3D humans under ideal conditions, requiring a clear and unobstructed scene. Those methods cannot be used to render humans in real-world scenes where obstacles may block the camera’s view and lead to partial occlusions. In this work, we present Wild2Avatar, a neural rendering approach catered for occluded in-the-wild monocular videos. We propose occlusion-aware scene parameterization for decoupling the scene into three parts - occlusion, human, and background. Additionally, extensive objective functions are designed to help enforce the decoupling of the human from both the occlusion and the background and to ensure the completeness of the human model. We verify the effectiveness of our approach with experiments on in-the-wild videos.
@@
[…]
1 Introduction
Rendering humans from videos has a wide variety of applications in many fields, including AR/VR [9, 12], film [3], and healthcare [6]. Videos from single cameras are widespread and easy to acquire, so rendering humans from monocular videos has been the target of copious research. Methods to accomplish this task such as Vid2Avatar [9], MonoHuman [51], and NeuMan [19] have achieved impressive performance. Despite being provided with only one camera view, these methods are able to render the human accurately from novel views. However, most existing human rendering methods are designed for ideal experimental scenes with little to no obstacles and a full view of the human in each frame. In real-world scenes, however, humans might move behind objects and move them around, causing the objects to block the camera’s view.
Most neural rendering methods face difficulty with real-world occluded scenarios for one main reason – lack of supervision. More specifically, in-the-wild scenarios usually do not provide ground-truth supervision of the appearance, shape, and pose of the human, making it necessary for the model to infer this information based on sparse evidence. This can be challenging, especially when a large portion of the human body is invisible. Due to the point-based rendering scheme of many neural methods, two very close coordinates can create a dramatic difference in rendering output when one is occluded and one is not [28, 46]. As a result, methods that are not adapted to occluded scenarios frequently display incomplete human bodies or render floaters and other artifacts when they encounter occlusions.
In this work, we address these shortcomings by introducing Wild2Avatar (Figure 1), an improved method for rendering humans under occlusion. Our method models the occlusions, human, and background as three separate neural fields, allowing a clean 3D reconstruction of the human regardless of occlusion. To accomplish this, we utilize scene self-decomposition. Extending upon inverted sphere parametrization [53], we propose occlusion-aware scene parametrization. In addition to the first sphere defined by inverted sphere parametrization, our parametrization introduces a second, inner sphere and defines the region from the camera to the edge of the inner sphere as the occlusion region. By rendering this region separately, we can decouple the occlusion from the rest of the scene. To ensure a high-fidelity and complete rendering of humans, we propose aggregating the three renderings through a combination of a pixel-wise photometric loss, a scene decomposition loss, an occlusion decoupling loss, and a geometry completeness loss. While there are past works designed to render occluded humans [26, 46], our work is the first to decouple the human from the occlusion cleanly, allowing for complete renderings of humans without floaters.
In summary, our contributions are: (i) We introduce occlusion-aware scene parametrization, a method to decouple a scene into three parts: occlusion, human body, and background. (ii) We propose a new rendering framework that renders each of these three parts separately, and design novel optimization objectives to ensure a clean decoupling of the occlusion and a more complete human rendering. (iii) We evaluate our method on challenging occlusion-intensive in-the-wild videos, demonstrating the effectiveness of our approach in rendering occluded humans.
@@ 2 Related Work
3D human modeling and rendering. There has been an abundance of recent work focused on free-viewpoint rendering of humans. While past works were able to achieve good quality renderings of humans from dense [10, 5, 54] and sparse [41, 55, 14, 47, 27, 25, 40] camera views, a recent research focus involves rendering a moving human from a single camera angle [44, 51, 9, 17, 18, 1, 2, 20, 11, 15]. To accomplish this task, Neural Radiance Fields (NeRFs) [32] based methods have been more favored recently due to their high rendering quality. Originally used to model static scenes, NeRFs have been adapted to model a dynamic human by parameterizing the human body using the SMPL [31] prior and mapping from the pose-independent canonical space to the observation space [44, 51, 9, 15]. HumanNeRF was one of the first to work on free-viewpoint rendering of humans in complex motion from monocular video [44]. MonoHuman models the deformation field with bi-directional constraints for even better rendering of humans [51]. While these methods perform well in rendering a high-fidelity human avatar from a monocular video in a clean environment, they fail when occlusions block the camera’s view of the human.
@@
Rendering via scene decomposition. Complex natural scenes are usually composed of multiple sub-components. So, a common approach to render these scenes is to model the different 3D components of the scene individually and then aggregate them together in the 2D image space [45, 19, 36, 9, 52]. GIRAFFE [36] was one of the pioneering works that proposed using generative neural fields to disentangle objects from the background. The same idea can be also found in NeRF++ [53], which is specifically designed for rendering unbounded surroundings, with two neural fields being trained separately for foreground and background. STNeRF [16] utilizes neural layers to render humans and background separately from multiple cameras, while NeuMan trains a separate human and scene radiance field on a monocular in-the-wild video [19]. Vid2Avatar uses inverse sphere parametrization from [53] to separate the dynamic human from the static background, resulting in clean and detailed human avatars [9]. Vid2Avatar’s high fidelity yet robust renderings with minimal artifacts inspire us to use it as the foundation of this work.
@@
Occlusion handling. While accounting for occlusions in rendering has been a long-standing research problem [48, 24, 35, 21, 33], occlusion handling for NeRFs is relatively new. OCC-NeRF uses a scene MLP and a background MLP to remove occlusions from the output [56]. NeuRay utilizes feature vectors to track the visibility of of each 3D point to determine whether it is occluded [29]. While these approaches reduce artifacts caused by occlusion, they require multiple input views to predict visibility, and are thus not generalizable to an in-the-wild monocular video of a dynamic human. HOSNeRF can render human, objects, and scene separately from a monocular in-the-wild video by introducing object bones that are attached to the human skeleton [26]. However, this approach does not account for object-based occlusions of the human. OccNeRF utilizes geometry and visibility priors with surface-based rendering to render occluded humans [46]. However, it renders numerous artifacts and floaters characteristic of NeRF-based methods (Figure 6). Our approach, on the other hand, renders the human body cleanly due to the separation of human, background, and occlusion.
@@
Figure 2: Wild2Avatar  renders occluded humans from a monocular video. It parameterizes the 3D scene into three parts: occlusion, human, and background, in order from closest to farthest from the camera. The human and the occlusions/background are individually modeled via separate neural radiance fields. The human is parameterized in a bounded sphere $\Pi$ by first deforming ray samples into a canonical space with the help of the pre-computed body pose $P$. The canonical points $\mathbf{x}$ are passed into a rendering network $\mathcal{F}^{fg}$ to learn the radiance $\mathbf{c}$ and distance $\mathbf{s}$ to the surface of the human, which can then be rendered through SDF-based volume rendering. The unbounded background is represented as coordinates on the surface of $\Pi$ along with their inverted distances. Another rendering network $\mathcal{F}^{scene}$ is used to learn the radiance and density for the background ray samples. The space of occlusion is determined as the interval between the camera and an inner sphere $\pi$. We parameterize ray samples as coordinates on the surface of $\pi$ and the negation of the inverted distances to the center of the inner sphere. We rely on the same network $\mathcal{F}^{scene}$ to render the occlusions. The three renderings are sequentially aggregated and supervised on a combination of losses, in which we specifically encourage the decoupling of the occlusion from the human through $\mathcal{L}_{occ}$ and penalize the incompleteness of human geometry through $\mathcal{L}_{comp}$.
@@ 3 Methods
In this section, we present Wild2Avatar, a model that renders 3D humans with complete geometry and high-fidelity appearance for in-the-wild monocular videos with occlusions. We start by reviewing the basics of neural radiance fields and the kinetics of the human body (section 3.1). We then present the key concepts of scene parameterization, which is central to the success of occlusion-aware human renderings (section 3.2). To follow, we outline the objective functions needed for complete human modeling and crisp scene decomposition (section 3.3). Lastly, we tie all the components together and outline the overall framework of Wild2Avatar in Figure 2.
@@ 3.1 Preliminaries and Background
Implicit neural radiance fields. Neural Radiance Fields (NeRFs) [32] learn a neural network $\mathcal{F}$ to model the mappings between (positionally embedded) 3D coordinates $\{\mathbf{x}\in\mathbb{R}^{3}\}$ and their radiance $\mathbf{c}(\mathbf{x})$ and density $\sigma(\mathbf{x})$. This representation is applicable to stationary one-object scenes but usually suffers from unexpected rendering artifacts and floaters in the wild. For rendering solid and continuous human bodies, using an implicit surface representation with a Signed Distance Function (SDF) is preferred [38, 49, 43, 9]. Instead of learning density values for each $\mathbf{x}$, $\mathcal{F}$ is trained to output distances to the surface of the human body $\mathbf{s}(\mathbf{x})$. The human geometry is then represented as a surface model by the zero-level set: $\{\mathbf{s}(\mathbf{x})=0\}$.
@@
where $\mathbf{B}$ is the bone transformation and $\{w^{i}_{(\cdot)}\}$ are vertex-wise skinning weights derived from SMPL [31] poses $\mathbf{P}$. With the above transformations, we are able to first optimize a static neural field for the dynamic human in the canonical space and then deform the ray samples to the observation space for volume rendering [39, 42, 40].
@@ 3.2 Occlusion-aware Scene Parameterization
To decouple the human from the background, the scene is usually parameterized separately for the human and background [9]. This kind of parameterization first introduced by [53] uses a sphere $\Pi$ to cover all of the space intended to be occupied by the human. The space outside the sphere is parameterized as the background by inverting the sphere. Rendering is then achieved via ray composition.
@@
Figure 3: Comparisons of scene parameterizations. (A) The regular two-part paradigm [53] parameterizes the foreground within a sphere $\Pi$ with radius $\mathbf{R}$ and background outside the sphere. (B) Our proposed occlusion-aware paradigm parameterizes the scene into three sequential parts. The occlusion is explicitly modeled within the interval space between the camera and an inner sphere $\pi$ with a radius $\mathbf{r}<\mathbf{R}$.
Human. For each frame in the video, we transform the deformed humans into the common canonical space, which is bounded by a sphere $\Pi$ with pre-defined radius $\mathbf{R}$. We learn a neural radiance field to model the surface of the human body implicitly (section 3.1). We follow [49, 9]’s approach to SDF-based volume rendering: for each canonical point $\mathbf{x}^{i}$, their signed distances $\mathbf{s}^{i}$ are first converted into density values $\sigma^{i}$ through scaled Laplace distribution’s Cumulative Distribution Function (CDF) to the negated signed distances. Then, we aggregate the radiance and density of each sample along the rays for volume rendering [30]:
@@
Background. The background of a scene includes the unbounded surroundings that cannot be naively parameterized by a sphere. We adopt the technique proposed by [53] that inverts the sphere $\Pi$ to represent every 3D point $\mathbf{x}=\{x,y,z\}$ as a quadruple $\mathbf{x}_{bg}=\{x_{bg},y_{bg},z_{bg},\frac{1}{d}\}$, where {$||\{x_{bg},y_{bg},z_{bg}\}||$ indicates 3D coordinates on the surface of $\Pi$ and $d$ is the Euclidean distance between the sphere origin $\mathbf{O}$ and $\mathbf{x}$. This inverted sphere parameterization enables sampling within the bounded interval $\frac{1}{d}\in(0,1]$. Given any $\frac{1}{d}$, $\{x_{bg},y_{bg},z_{bg}\}$ can be calculated by rotating the vector [53]. Similar to [9], the rendering of the background is conditioned on the combination of a per-frame latent code, 4D parameterization, and embedded ray directions.
Occlusion. On top of the two-part scene parameterization, in this work, we propose to cater for obstacles that come between the human and the camera when filming the video. We define the occlusion to lie inside the sphere $\Pi$ and occupy a bounded sub-space within $\Pi$. To parameterize the human and occlusion separately, a concentric inner sphere $\pi$ is introduced with radius $\mathbf{r}<\mathbf{R}$. Without any semantic priors about the obstacles (e.g. size, appearance), we use the entire space between the camera and $\pi$ to learn the neural radiance field for occlusion. To fully bound this sub-space, $\pi$ is built as an inscribed sphere to the outermost rays shot from the camera, and its radius $\mathbf{r}$ is determined accordingly to ensure every ray intersects $\pi$ at least once. For any radius $r$, the number of intersections can be determined via:
$(\mathbf{o}\cdot\mathbf{d})^{2}-||\mathbf{o}^{2}||+r^{2},$ (4)
where $\mathbf{o}$ is the camera location, $\mathbf{d}$ is the ray direction, and |||||$\cdot$||| | ⋅ | | is L1 norm on the coordinate dimension. There is no intersection between the rays and $\pi$ when the above function evaluates to negative, and there must be at least one intersection otherwise. Given the monotonicity of the function w.r.t $r$, we can easily find the minimum possible $\mathbf{r}$ through binary search within $(0,\mathbf{R})$ during pre-processing.
Since we do not focus on rendering quality for the obstacles and background, we improve network efficiency by using the same rendering network $\mathcal{F}^{Scene}$ for both obstacles and background while conditioning on different per-frame latent codes. For the unbounded background, the quadruple $\mathbf{x}_{bg}$ must be computed through vector rotation which is time-consuming. In the bounded occlusion space, the quadruple $\mathbf{x}_{occ}$ can be computed more efficiently by solving the following quadratic equation w.r.t $t$:
$\begin{gathered}||\mathbf{o}+t\mathbf{\mathbf{d}}||=\mathbf{r},\\ ||\mathbf{d}||^{2}t^{2}+2(\mathbf{o}\cdot\mathbf{d})t+||\mathbf{o}||^{2}-\mathbf{r}=\vec{0},\\ t=\frac{-\sqrt{||\mathbf{\mathbf{o}}\cdot\mathbf{d}||^{2}+\left(||\mathbf{r}^{2}||-||\mathbf{\mathbf{o}}^{2}||\right)||\mathbf{d}^{2}||}-||\mathbf{\mathbf{o}}\cdot\mathbf{d}||}{||\mathbf{d}^{2}||}.\end{gathered}$ (5)
@@
Sharing the same network $\mathcal{F}^{scene}$ for both background and occlusion requires the inputs to be at the same magnification scale. To do this, we normalize both $\{x_{occ},y_{occ},z_{occ}\}$ and $\{x_{bg},y_{bg},z_{bg}\}$ to the unit sphere. Such normalization will not lead to overlapped surface vectors since $\mathbf{r}$ is strictly smaller than $\mathbf{R}$. To further decouple the dependencies for occlusion/background rendering, we determine the quadruple $\mathbf{x}_{occ}$ as the concatenation of $\{x_{occ},y_{occ},z_{occ}\}$ and negation of the inverted depth $-\frac{\mathbf{r}}{||\mathbf{x}^{2}||}$.

### v2

# Rendering Humans Behind Occlusions

@@ Abstract
Rendering the visual appearance of moving humans from occluded monocular videos is a challenging task. Most existing research renders 3D humans under ideal conditions, requiring a clear and unobstructed scene. Those previous methods cannot be used to render humans in real-world scenes where obstacles may block the camera’s view and lead to partial occlusions. In this work, we present Wild2Avatar, a neural rendering approach catered for occluded in-the-wild monocular videos. We propose occlusion-aware scene parameterization for decoupling the scene into three parts - occlusion, human, and background. Additionally, extensive objective functions are designed to help enforce the decoupling of the human from both the occlusion and the background and to ensure the completeness of the human model. Wild2Avatar is verified with experiments on 14 challenging in-the-wild videos.
Index Terms:
Human modeling, Neural rendering, Occlusion handling.
1 Introduction
Rendering humans from videos has a wide variety of applications in many fields, including augmented/virtual reality [1, 2], film [3], and healthcare [4]. Videos from single cameras are widespread and easy to acquire, so rendering humans from monocular videos has been the target of copious research. Methods to accomplish this task such as Vid2Avatar [1], MonoHuman [5], and NeuMan [6] have achieved impressive performance. However, most existing human rendering methods are designed for ideal experimental scenes with little to no obstacles and a full view of the human in each frame. In real-world scenes, however, humans might move behind objects and move them around, causing the objects to block the camera’s view.
Most neural rendering methods face significant challenges when dealing with occlusions in real-world scenarios, primarily due to insufficient supervision. Unlike controlled environments, in-the-wild settings typically lack ground-truth reference for human appearance, geometry, and pose. As a result, the model must infer these attributes from limited and often partial visual evidence. This inference becomes especially difficult when large portions of the human body are occluded. Additionally, many existing methods rely on point-based rendering schemes, which can produce vastly different outputs when adjacent points are occluded [7, 8]. Methods that are not specifically adapted to handle occlusions tend to generate incomplete human bodies with visual artifacts and floaters. This problem is particularly critical in surveillance videos from environments such as construction sites and hospitals, where humans are frequently blocked by various objects. Accurate reconstruction and rendering of occluded humans in these contexts is therefore essential, as it directly affects performance in downstream tasks.
@@ 1 Introduction
In this work, we address these shortcomings by introducing Wild2Avatar (Figure 1), an improved method for rendering higher-fidelity humans under occlusion based on our previous ICCV-2023 publication OccNeRF [8].
Our method builds off a novel scene parameterization technique, occlusion-aware scene parametrization that divides the whole scene into three parts — occlusions, human, and background. By extending upon inverted sphere parametrization [9] with a second, inner sphere, and by defining the region from the camera to the edge of the inner sphere as the occlusion region, our method allows for a clean 3D reconstruction of the human regardless of occlusion. To ensure a high-fidelity and complete rendering of humans, we propose aggregating the three components through a combination of a pixel-wise photometric loss, a scene decomposition loss, an occlusion decoupling loss, and a geometry completeness loss. Along with OccNeRF [8], our work is the first neural method for cleanly rendering humans from occluded monocular videos free of floaters. Similar to most previous literature in human rendering [10, 1, 8], our method mainly focuses on reconstructing the human – we do not aim to accurately render irrelevant occlusions or the background in this work.
In summary, our contributions in this paper are:
  1. 1.
    We introduce occlusion-aware scene parametrization, a method to decouple a scene into three parts: occlusion, human body, and background.
  2. 2.
    We propose a new rendering framework that renders each of these three parts separately and design novel optimization objectives to ensure a clean decoupling of the occlusion and a complete human rendering.
  3. 3.
    We evaluate our method on 14 challenging occlusion-intensive in-the-wild videos, demonstrating its effectiveness in rendering occluded humans.
@@ 2 Related Work
There have been numerous works focused on free-viewpoint rendering of humans. While past works were able to achieve good quality renderings of humans from dense [11, 12, 13] and sparse [14, 15, 16, 17, 18, 19, 20] camera views, a recent research focus involves rendering a moving human from a single camera angle [10, 5, 1, 21, 22, 23, 24, 25, 26, 27]. To accomplish this task, Neural Radiance Fields (NeRFs) [28] based methods have been more favored recently due to their high rendering quality. Originally used to model static scenes, NeRFs have been adapted to model a dynamic human by parameterizing the human body using the SMPL [29] prior and mapping from the pose-independent canonical space to the observation space [10, 5, 1, 27]. HumanNeRF was one of the first to work on free-viewpoint rendering of humans in complex motion from monocular video [10]. MonoHuman models the deformation field with bi-directional constraints for better rendering of humans [5].
More recent works in neural rendering harness 3D Gaussian splatting [30], which allows for extremely efficient training and rendering of scenes. By initializing the Gaussians based on the SMPL prior, numerous concurrent works have been able to render humans extremely efficiently and accurately [31, 32, 33, 34, 35].
However, while these existing methods perform well in a clean environment, they mostly fail when occlusions block the camera’s view of the human.
2.2 Rendering via Scene Decomposition
@@
Complex natural scenes are usually composed of multiple sub-components. So, a common approach to render these scenes is to model the different 3D components of the scene individually and then aggregate them together in the 2D image space [36, 6, 37, 1, 38]. GIRAFFE [37] was one of the pioneering works that proposed using generative neural fields to disentangle objects from the background. The same idea can be also found in NeRF++ [9], which is specifically designed for rendering unbounded surroundings, with two neural fields being trained separately for foreground and background. STNeRF [39] utilizes neural layers to render humans and background separately from multiple cameras, while NeuMan [6] trains a separate human and scene radiance field on a monocular in-the-wild video. In section 4.3, we detail why NeuMan is not able to be used as a major comparison baseline. Vid2Avatar uses inverse sphere parametrization from [9] to separate the dynamic human from the static background, resulting in clean and detailed human avatars [1]. Vid2Avatar’s high-fidelity yet robust renderings with minimal artifacts inspire us to use it as the foundation of this work. More recently, HSR [40] introduced a unified framework that jointly reconstructs dynamic humans and static scenes in a shared global coordinate system from monocular RGB videos. Unlike prior works that process humans and scenes separately, HSR explicitly models physical interactions, such as occlusions and interpenetration, enabling more coherent and realistic reconstructions.
2.3 Occlusion Handling
@@
While accounting for occlusions in rendering has been a long-standing research problem [41, 42, 43, 44, 45], occlusion handling for NeRFs is relatively new. OCC-NeRF [46] uses a scene MLP and a background MLP to remove occlusions from the output. NeuRay [47] utilizes feature vectors to track the visibility of each 3D point to determine whether it is occluded. While these approaches reduce artifacts caused by occlusion, they require multiple input views to predict visibility and are thus not generalizable to an in-the-wild monocular video of a dynamic human. HOSNeRF [48] can render human, objects, and the scene separately from a monocular in-the-wild video by introducing object bones that are attached to the human skeleton. However, this approach is only robust to very small-scale occlusions and is only applicable for dynamic scenes due to the constraints of COLMAP. We detail why HOSNeRF is not able to be used as a major comparison baseline in section 4.3.
2.4 Connection to Our Previous Work
We first introduced the problem of occluded human rendering from monocular videos along with a solution in our previous ICCV-2023 publication OccNeRF [8]. This paper is a significant extension with the following three major improvements:
  1. 1.
    We have thoroughly revised the OccNeRF pipeline and present novel methodologies for both scene modeling and human rendering.
  2. 2.
    We have evaluated the proposed methods on seven more challenging in-the-wild videos with real-world occlusions. Extensive experiments were conducted for further study on the robustness towards the number of training images as well as the model’s generative ability.
  3. 3.
    We achieve superior results that renders high-fidelity human bodies with much less artifacts and floaters.
@@ 2.4 Connection to Our Previous Work
Figure 2: Wild2Avatar  renders occluded humans from a monocular video. It parameterizes the 3D scene into three parts: occlusion, human, and background, in order from closest to farthest from the camera. The human and the occlusions/background are individually modeled via separate neural radiance fields. The human is parameterized in a bounded space between an inner sphere $\Pi$ with radius $\mathbf{r}$ and an outer sphere with radius $\mathbf{R}$ by first deforming ray samples into a canonical space with the help of the pre-computed body pose $P$. The canonical points $\mathbf{x}$ are passed into a rendering network $\mathcal{F}^{fg}$ to learn the radiance $\mathbf{c}$ and distance $\mathbf{s}$ to the surface of the human, which can then be rendered through SDF-based volume rendering. The unbounded background samples outside the outer sphere are represented as coordinates on the surface of the outer sphere along with their inverted distances. Another rendering network $\mathcal{F}^{scene}$ is used to learn the radiance and density for the background ray samples (shown as green dots ). The space of occlusion is determined as the interval between the camera and the inner sphere $\pi$. We parameterize ray samples as coordinates on the surface of $\pi$ and the negation of the inverted distances to the center of the inner sphere. We rely on the same network $\mathcal{F}^{scene}$ to render the occlusions. The three renderings are sequentially aggregated and supervised on a combination of losses, in which we specifically encourage the decoupling of the occlusion from the human through $\mathcal{L}_{occ}$ and penalize the incompleteness of human geometry through $\mathcal{L}_{comp}$.
@@ 3 Methods
We present Wild2Avatar, a model that renders 3D humans with complete geometry and high-fidelity appearance for in-the-wild monocular videos with occlusions. We start by reviewing the basics of NeRFs and the kinetics of the human body (section 3.1). We then present the key concepts of scene parameterization, which is central to the success of occlusion-aware human renderings (section 3.2). To follow, we outline the objective functions needed for complete human modeling and crisp scene decomposition (section 3.3). Lastly, we tie all the components together and outline the overall framework of Wild2Avatar in Figure 2.
@@ 3.1 Preliminaries and Background
Implicit neural radiance fields. NeRFs [28] learn a neural network $\mathcal{F}$ to model the mappings between (positionally embedded) 3D coordinates $\{\mathbf{x}\in\mathbb{R}^{3}\}$ and their radiance $\mathbf{c}(\mathbf{x})$ and density $\sigma(\mathbf{x})$. This representation is applicable to stationary one-object scenes but usually suffers from unexpected rendering artifacts and floaters in the wild. For rendering solid and continuous human bodies, using an implicit surface representation with a Signed Distance Function (SDF) is preferred [49, 50, 51, 1]. Instead of learning density values for each $\mathbf{x}$, $\mathcal{F}$ is trained to output distances to the surface of the human body $\mathbf{s}(\mathbf{x})$. The human geometry is then represented as a surface model by the zero-level set: $\{\mathbf{s}(\mathbf{x})=0\}$.
@@
where $\mathbf{B}$ is the bone transformation and $\{w^{i}_{(\cdot)}\}$ are vertex-wise skinning weights derived from SMPL [29] poses $\mathbf{P}$. We are then able to first optimize a static neural field for the dynamic human in the canonical space and then deform the ray samples to the observation space for volume rendering [53, 54, 20].
@@ 3.2 Occlusion-aware Scene Parameterization
To decouple the human from the background, the scene is usually parameterized separately [1]. This kind of parameterization first introduced by [9] uses a sphere $\Pi$ to cover all of the space intended to be occupied by the human. The space outside the sphere is parameterized as the background by inverting the sphere (Figure 3 (A)). Rendering is then achieved via ray composition.
@@
Figure 3: Different Scene parameterizations in 2D simplifications. (A) The regular two-part paradigm [9] parameterizes the foreground within a sphere $\Pi$ with radius $\mathbf{R}$ and background outside the sphere. (B) Our proposed occlusion-aware paradigm parameterizes the scene into three sequential parts. The occlusion is explicitly modeled within the interval space between the camera and an inner sphere $\pi$ with a radius $\mathbf{r}<\mathbf{R}$.
Human. For each frame in the video, we transform the deformed humans into the common canonical space, which is bounded by a sphere $\Pi$ with pre-defined radius $\mathbf{R}$. We learn a neural radiance field to model the surface of the human body implicitly (section 3.1). We follow [50, 1]’s approach to SDF-based volume rendering. For each point $\mathbf{x}^{i}$ in the canonical space, we first compute its signed distance $\mathbf{s}^{i}$ and then transform signed distances into density values $\sigma^{i}$ by applying the cumulative distribution function (CDF) of the scaled Laplace distribution to the negative of the signed distances. Then, we aggregate the radiance and density of each sample along the rays for volume rendering [55]:
@@
Background. To handle the unbounded environment outside a scene’s foreground, we adopt the inverted-sphere scheme from NeRF++ [9]. We treat any point whose distance from the origin exceeds the radius $\mathbf{R}$ of an outer sphere $\Pi$ as background. Such a point $\mathbf{x}=[x,y,z]$ is represented as the 4-tuple $\mathbf{x}_{\text{bg}}=[x_{\text{bg}},y_{\text{bg}},z_{\text{bg}},1/d]$: the first three entries store its direction on the outer sphere surface, while $1/d$ (inverse distance) compresses the infinite radial range into the bounded interval $(0,1]$. Sampling uniformly in $1/d$ concentrates rays near the foreground, preserves coverage of far-field content, and keeps all coordinates numerically stable [9]. Similar to [1], the rendering of the background is conditioned on the combination of a per-frame latent code that is constantly optimized during training, 4D parameterization, and embedded ray directions.
[…]
Since we do not focus on rendering quality for the obstacles and background, we improve network efficiency by using the same rendering network $\mathcal{F}^{Scene}$ for both obstacles and background while conditioning on different per-frame latent codes. For the unbounded background, the quadruple $\mathbf{x}_{bg}$ must be computed through vector rotation which is time-consuming. In the bounded occlusion space, the quadruple $\mathbf{x}_{occ}$ can be computed more efficiently by solving the following quadratic equation with respect to $t$:
[…]
@@
[…]